# Spatial-Temporal Decoupling Contrastive Learning for Skeleton-based Human Action Recognition

### v3

@@ 3.2.1 Spatial Temporal Feature Decoupling
The details of the proposed STFD are shown in (b) of Figure 3. The feature $\boldsymbol{X}$ is fed into two parallel branches for feature decoupling. In the spatial feature decoupling (SFD) module, a temporal pooling layer is adopted to obtain the spatial attentive feature $\boldsymbol{X}_{s}\in\mathbb{R}^{J\times C}$. And the temporal attentive feature $\boldsymbol{X}_{t}\in\mathbb{R}^{T\times C}$ is obtained through a spatial pooling layer. Then, two linear transformation functions $\theta$ and $\phi$ with a channel reduction rate of $r$ are used to transform the attentive features into a neatly compact representation as follows:
$\left\{\begin{matrix}\boldsymbol{X}_{s}^{*}=\theta(\boldsymbol{X}_{s})=\boldsymbol{X}_{s}\boldsymbol{W}_{\theta}\in\mathbb{R}^{J\times C/r}\\ \\ \boldsymbol{X}_{t}^{*}=\phi(\boldsymbol{X}_{t})=\boldsymbol{X}_{t}\boldsymbol{W}_{\phi}\in\mathbb{R}^{T\times C/r}\end{matrix}\right.$ (4)
@@
where, $\boldsymbol{W}_{\theta},\boldsymbol{W}_{\phi}\in R^{C\times C/r}$ are the weights of the spatial and temporal transformation function, respectively. After this, to retain the attentive information, we squeeze the spatial dimension into the channel dimension in the SFD module, and the TFD module is similar. In the end, two liner transformation functions $\zeta$ and $\sigma$ are used to transform the features to the latent space for further contrastive learning, which can be formulated as:
$\left\{\begin{matrix}\boldsymbol{s}=\zeta(\boldsymbol{s}^{*})=\boldsymbol{s}^{*}\boldsymbol{W}_{\zeta}\in\mathbb{R}^{D}\\ \\ \boldsymbol{t}=\sigma(\boldsymbol{t}^{*})=\boldsymbol{t}^{*}\boldsymbol{W}_{\sigma}\in\mathbb{R}^{D}\end{matrix}\right.$ (5)
@@
where, $\boldsymbol{W}_{\zeta}\in\mathbb{R}^{JC/r\times D}$, $\boldsymbol{W}_{\sigma}\in\mathbb{R}^{TC/r\times D}$ are the weights of final projection function. In this way, we decouple the spatial attentive feature $\boldsymbol{s}$ and the temporal attentive feature $\boldsymbol{t}$ from the spatial-temporal feature $\boldsymbol{X}$.
@@ 3.2.2 Memory Bank
To obtain abundant negative pairs, we contrast two memory banks $\mathcal{M}_{spa}$ and $\mathcal{M}_{tem}$ to store the decoupled attentive features. Specifically, to avoid introducing hyperparameters and encourage the model to explore the cross-sequence context fully, we set the two memory banks to store the attentive features of all samples. Thus, the memory banks $\mathcal{M}_{spa},\mathcal{M}_{tem}\in\mathbb{R}^{L\times D}$. $L$ is the number of the instance of the dataset. When an instance is fed into the model to extract features, the stored features in the memory banks are updated according to the corresponding index.
$\left\{\begin{matrix}\mathcal{M}_{spa}[i]=\boldsymbol{s}_{i}\\ \\ \mathcal{M}_{tem}[i]=\boldsymbol{t}_{i}\end{matrix}\right.$ (6)
where $i$ is the index of the instance in the dataset. Each element in $\mathcal{M}_{spa}$ and $\mathcal{M}_{tem}$ indicates a decoupled feature embedding from an instance. In the memory banks, the corresponding labels with the decoupled features are stored for the feature selection. Particularly, the decoupled features with the same category label as the current instance are selected as positive pairs, and negative pairs are sampled from the memory banks with other labels.
@@ 3.2.3 Training Objective
The positive pairs and negative pairs sampled from the two memory banks are used to achieve spatial-temporal decoupled contrastive learning. To pull together the positive pairs and push away the negative pairs, we define the distances between the two vectors with cosine similarity as follows:
$\left\langle\boldsymbol{u},\boldsymbol{v}\right\rangle=\frac{\boldsymbol{u}\cdot\boldsymbol{v}^{T}}{||\boldsymbol{u}||\;||\boldsymbol{v}||}$ (7)
where $\boldsymbol{u}$, $\boldsymbol{v}$ are the features sampled from the memory banks. Then the InfoNCE loss adopted for spatial-specific feature contrasting and temporal-specific feature contrasting can be written as follows:
$\mathcal{L}_{NCE}^{spa}=-\sum\limits_{\boldsymbol{s}^{+}\in\mathcal{M}^{+}_{spa}}log\frac{\left\langle\boldsymbol{s},\boldsymbol{s}^{+}\right\rangle/\tau}{\left\langle\boldsymbol{s},\boldsymbol{s}^{+}\right\rangle/\tau+\sum\limits_{\boldsymbol{s}^{-}\in\mathcal{M}^{-}_{spa}}\left\langle\boldsymbol{s},\boldsymbol{s}^{-}\right\rangle/\tau}$
$\mathcal{L}_{NCE}^{tem}=-\sum\limits_{\boldsymbol{t}^{+}\in\mathcal{M}^{+}_{tem}}log\frac{\left\langle\boldsymbol{t},\boldsymbol{t}^{+}\right\rangle/\tau}{\left\langle\boldsymbol{t},\boldsymbol{t}^{+}\right\rangle/\tau+\sum\limits_{\boldsymbol{t}^{-}\in\mathcal{M}^{-}_{tem}}\left\langle\boldsymbol{t},\boldsymbol{t}^{-}\right\rangle/\tau}$
where $\tau$ is a temperature hyperparameter in contrastive learning. $\mathcal{M}^{+}_{spa}$ and $\mathcal{M}^{+}_{tem}$ are the sets of the decoupled features from the positive samples with the same label as the current feature in the memory banks. $\mathcal{M}^{-}_{spa}$ and $\mathcal{M}^{-}_{tem}$ are the sets of negative features with different lalbes.
Finally, the total training loss function of our proposed STD-CL is defined as follows:
@@ 3.2.4 Sampling Strategy
To achieve efficient contrast across sequences, the strategy of sampling features from memory banks for contrastive learning has a large impact on the experimental performance. Therefore, in this paper, we follow the sampling strategy in Huang et al. (2023), which combines the hard mining strategy and random sampling strategy. The hard mining strategy is introduced to focus on the hard samples, which means the $N^{+}_{H}$ hardest positive samples with the lowest similarity and $N^{-}_{H}$ hardest negative samples with the highest similarity in equation 7. Furthermore, the random sampling strategy is adopted to maintain global random contrast. There are $N^{-}_{R}$ random negative samples selected to contrast with the input instance. In total, the sampling strategy adopted in our STD-CL takes contrast efficiency into account and contributes to the hard examples.
@@ 4 Experiment
AGC-LSTM Si et al. (2019) - - 93.3* -
DC-GCN+ADG Cheng et al. (2020a) - - 95.3 -
Shift-GCN Cheng et al. (2020b) 92.5 - 94.2 94.6
2S-AGCN Shi et al. (2019) 92.0 92.2 95.0 95.5
2S-AGCN w/STD-CL 93.8 93.1 95.9 97.0
CTR-GCN Chen et al. (2021a) 94.6 91.8 94.2 96.5
CTR-GCN w/STD-CL 94.8 94.6 95.7 97.2
InfoGCN Chi et al. (2022) 93.8 94.2 95.5 96.1
Info-GCN w/STD-CL 94.2 95.5 95.7 96.3
Hyperformer Zhou et al. (2022) 92.7 95.0 95.0 96.6
Hyperformer w/STD-CL 95.3 95.3 96.3 97.0
4.1 Dataset Settings
NTU RGB+D. NTU60 Shahroudy et al. (2016) is the most widely used dataset for skeleton-based action recognition. It contains 56,880 samples in 60 action classes. These samples are performed by 40 distinct subjects. It has been configured into two benchmarking structures: cross-subject (X-sub) and cross-view (X-view). In the X-sub, sequences from 20 individuals are allocated for training purposes, with the remaining 20 sets of sequences earmarked for validation. Meanwhile, under the X-view, skeleton sequences are divided based on camera views; two camera views are used for training while the remaining ones are utilized for evaluation.
NTU RGB+D 120. NTU120 dataset Liu et al. (2019) adds 57367 new skeleton sequences and 60 new action classes to the original NTU60 dataset. There are 32 various configurations in it, each of which depicts a different location and background. The authors offered the cross-subject (X-sub) and cross-setup (X-set) as two benchmark evaluations. In the X-sub, sequences from 53 subjects are for training, and the other 53 subjects are for testing. In the X-set, skeleton sequences are split by setup ID. Samples from even set-up IDs are used for training, and the odds are used for evaluation.
Northwestern-UCLA. NW-UCLA dataset Wang et al. (2014) contains 1494 video clips of 10 different actions captured from three Kinect cameras. We follow the same evaluation protocol in Wang et al. (2014): the first two cameras for training and the other for testing.
@@ 4.1 Dataset Settings
Baseline Chen et al. (2021a) 84.9
w/GAP Xiang et al. (2023) 85.5
w/SkeletonGCL Huang et al. (2023) 85.6
w/FR Head Zhou et al. (2023) 85.5
w/STD-CL 85.6
STD-CL(Hyperformer Zhou et al. (2022)) 86.8
4.2 Implementation Details
To fully validate the effectiveness and generalizability of STD-CL, we take four different model-based approaches as baseline models. (1) CNN-based method. We select HCN Li et al. (2018) as the CNN-based method to validate STD-CL. We reproduce the experimental results using the released code and conduct our STD-CL on it. And we follow the training recipes as the paper describes. (2) GCN-based method. 2S-AGCN Shi et al. (2019) and CTR-GCN Chen et al. (2021a), the widely used GCN-based model for skeleton-based action recognition. For CTR-GCN, we follow their training strategies. Particularly, for 2S-AGCN, we update the training strategies from CTR-GCN, which can improve the performance significantly. (3) Transformer-based method. For transformer-based methods, we choose Hyperformer as the baseline. More implementation details are presented in the supplementary material.
4.3 Comparsion With The State-of-The-Art
We compare our method with previous SOTA methods in Tables 1 and 2. To fully validate the proposed STD-CL, we report the experimental results in a single modality and the ensemble results of multi-modality. $J$ means the joint stream, $B$ denotes the bone stream, $2S$ denotes the ensemble results of the above two streams in default, and $4S$ incidents the ensemble results of additional joint motion and bone motion streams. Particularly, due to the design of their model architecture, the experimental results of VA-LSTM Zhang et al. (2017), ACG-LSTM Si et al. (2019) and HCN Li et al. (2018) are the ensemble results of two streams. Therefore, we classify the results into $2S$.
As shown in Table 1, on NTU60 and NTU120, combined with STD-CL, all four various baselines achieve solid gains on these four benchmarks of the two datasets over different settings. Moreover, the improvement in a single stream can be generalized to the results of the multi-stream ensemble. Take CTR-AGCN as an example, it improves by 0.6% (89.8% to 90.4%) in the joint stream, 0.6% (90.2% to 90.8%) in the bone stream, and 0.2% (92.0% to 92.2%) in their fusion result.
As shown in Table 2, on NW-UCLA, STD-CL outperforms 2S-AGCN by 1.5%. It also outperforms CTR-CCN by 0.7%. And it outperforms the recent work InfoGCN and Hyperformer by 0.2% and 0.4%, respectively. Considering that the performance on this dataset is already very high, such improvement brought by STD-CL is significant.
4.4 Comparsion With Other Contrastive Learning Methods

### v4

@@ 3.2 Spatial Temporal Feature Decomposition
The details of the proposed STFD are shown in Figure 3. The representation $\bm{X}$ is fed into two parallel branches for feature decomposition. Taking the spatial feature decomposition (SFD) module as an example, a temporal pooling layer is adopted to obtain the spatial attentive feature $\bm{X}_{s}\in\mathbb{R}^{J\times C}$. And the temporal-specific feature $\bm{X}_{t}\in\mathbb{R}^{T\times C}$ is obtained through a spatial pooling layer. Then, two linear transformation functions $\theta$ and $\phi$ with a channel reduction rate of $r$ are used to transform the attentive features into neatly compact representations as follows:
$\left\{\begin{matrix}\bm{X}_{s}^{*}=\theta(\bm{X}_{s})=\bm{X}_{s}\bm{W}_{\theta}\in\mathbb{R}^{J\times C/r}\\ \\ \bm{X}_{t}^{*}=\phi(\bm{X}_{t})=\bm{X}_{t}\bm{W}_{\phi}\in\mathbb{R}^{T\times C/r}\end{matrix}\right.$ (4)
@@
where $\bm{W}_{\theta},\bm{W}_{\phi}\in R^{C\times C/r}$ are the weights of the spatial and temporal transformation function, respectively. After this, to retain the attentive information, we squeeze the spatial dimension into the channel dimension in the SFD module, and the TFD module is similar. In the end, two liner transformation functions $\zeta$ and $\sigma$ are used to transform the features to the latent space for further contrastive learning, which can be formulated as:
$\left\{\begin{matrix}\bm{s}=\zeta(\bm{s}^{*})=\bm{s}^{*}\bm{W}_{\zeta}\in\mathbb{R}^{D}\\ \\ \bm{t}=\sigma(\bm{t}^{*})=\bm{t}^{*}\bm{W}_{\sigma}\in\mathbb{R}^{D}\end{matrix}\right.$ (5)
@@
where, $\bm{W}_{\zeta}\in\mathbb{R}^{JC/r\times D}$, $\bm{W}_{\sigma}\in\mathbb{R}^{TC/r\times D}$ are the weights of final projection function. In this way, we decompose the spatial-specific feature $\bm{s}$ and the temporal-specific feature $\bm{t}$ from the spatiotemporal representation $\bm{X}$.
Thus, the spatiotemporal representation extracted from the skeleton encoder is decomposed into two compact and specific features along the respective dimensions.
3.3 Contrastive Spatiotemporal Representation Enhancement
Figure 4: Framework of the proposed CSRE. The black dotted lines represent the training process, and the red solid line represents the testing process. The skeleton encoder extracts spatial-temporal representation $\bm{X}$ from the input skeleton sequence $\bm{S}$. Then,
the $\bm{X}$ is fed into the STFD module to decompose spatial attentive feature $\bm{s}$ and temporal attentive feature $\bm{t}$. Two memory banks are built to store the decomposed features respectively. We contrast the positive and negative instance features sampled from the memory banks according to the labels. The memory banks and the contrastive learning process are only used in the training phase and have no impact in the testing phase.
The framework of the proposed CSRE is shown in Figure 4. To further enhance the distinctiveness of the representation $\bm{X}$, helping distinguish ambiguous samples, we apply the decomposed components to contrastive learning to explore the cross-sequence semantic relations. The core is to explore the latent data distributions by pulling together the features from the positive pairs and pushing away the negative pairs.
@@ 3.3.1 Memory Bank
To obtain abundant negative pairs, we construct two memory banks $\mathcal{M}_{spa}$ and $\mathcal{M}_{tem}$ to store the decomposed features. Specifically, to avoid introducing hyperparameters and fully encourage the model to explore the cross-sequence context, we set the two memory banks to store the attentive features of all samples. Thus, the memory bank $\mathcal{M}_{spa},\mathcal{M}_{tem}\in\mathbb{R}^{L\times D}$. $L$ is the number of the instance of the dataset. When an instance is fed into the network to extract features, the stored features in the memory banks are updated according to the corresponding index.
$\left\{\begin{matrix}\mathcal{M}_{spa}[i]=\bm{s}_{i}\\ \\ \mathcal{M}_{tem}[i]=\bm{t}_{i}\end{matrix}\right.$ (6)
where $i$ is the index of the instance in the dataset. Each element in $\mathcal{M}_{spa}$ and $\mathcal{M}_{tem}$ indicates a decomposed feature embedding from an instance. In the memory banks, the corresponding labels with the decomposed features are stored for the feature selection. Particularly, the specific features with the same category label as the current instance are selected as positive pairs, and negative pairs are sampled from the memory banks with other labels.
@@ 3.3.2 Training Objective
The positive pairs and negative pairs sampled from the two memory banks are used to achieve contrastive learning. To pull together the positive pairs and push away the negative pairs effectively, we define the distances between the two vectors with cosine similarity as follows:
$\left\langle\bm{u},\bm{v}\right\rangle=\frac{\bm{u}\cdot\bm{v}^{T}}{||\bm{u}||\;||\bm{v}||}$ (7)
where $\bm{u}$, $\bm{v}$ are the features sampled from the memory banks. Then the InfoNCE loss adopted for spatial-specific feature contrasting and temporal-specific feature contrasting can be written as follows:
$\mathcal{L}_{NCE}^{spa}=-\sum\limits_{\bm{s}^{+}\in\mathcal{M}^{+}_{spa}}log\frac{\left\langle\bm{s},\bm{s}^{+}\right\rangle/\tau}{\left\langle\bm{s},\bm{s}^{+}\right\rangle/\tau+\sum\limits_{\bm{s}^{-}\in\mathcal{M}^{-}_{spa}}\left\langle\bm{s},\bm{s}^{-}\right\rangle/\tau}$ (8)
$\mathcal{L}_{NCE}^{tem}=-\sum\limits_{\bm{t}^{+}\in\mathcal{M}^{+}_{tem}}log\frac{\left\langle\bm{t},\bm{t}^{+}\right\rangle/\tau}{\left\langle\bm{t},\bm{t}^{+}\right\rangle/\tau+\sum\limits_{\bm{t}^{-}\in\mathcal{M}^{-}_{tem}}\left\langle\bm{t},\bm{t}^{-}\right\rangle/\tau}$ (9)
where $\tau$ is a temperature hyperparameter in contrastive learning. $\mathcal{M}^{+}_{spa}$ and $\mathcal{M}^{+}_{tem}$ are the sets of the attentive features from the positive samples with the same label as the current feature in the memory banks. $\mathcal{M}^{-}_{spa}$ and $\mathcal{M}^{-}_{tem}$ are the sets of negative features with different lalbes.
Finally, the total training loss function of our proposed CSRE is defined as follows:
@@ 3.3.3 Sampling Strategy
To achieve efficient contrast across sequences, the strategy of sampling features from memory banks for contrastive learning has a large impact on the experimental performance. Therefore, in this paper, we follow the hybrid sampling strategy in [33], which combines the hard mining strategy and random sampling strategy. The hard mining strategy is introduced to focus on the hard samples, which means the $N^{+}_{H}$ hardest positive samples with the lowest similarity and $N^{-}_{H}$ hardest negative samples with the highest similarity in equation 7. Furthermore, the random sampling strategy is adopted to maintain global random contrast. There are $N^{-}_{R}$ random negative samples selected to contrast with the input instance. In total, the sampling strategy adopted in our CSRE takes contrast efficiency into account and contributes to the hard examples. More details about the sampling strategy are in Section 4.5
@@ 4 Experiment
4.1 Dataset Settings
NTU RGB+D. NTU60 [35] is the most widely used dataset for skeleton-based action recognition. It contains 56,880 samples in 60 action classes. These samples are performed by 40 distinct subjects. It has been configured into two benchmarking structures: cross-subject (X-sub) and cross-view (X-view). In the X-sub, sequences from 20 individuals are allocated for training purposes, with the remaining 20 sets of sequences earmarked for validation. Meanwhile, under the X-view, skeleton sequences are divided based on camera views; two camera views are used for training, while the remaining ones are utilized for evaluation.
[…]
Dataset NTU 60
Setting X-Sub X-View
Method/Streams J B 2S 4S J B 2S 4S
ST-GCN [16] 81.5 - - - 88.3 -
Shift-GCN [36] 87.8 - 89.7 90.7 95.1 - 96.0 96.5
TA-CNN [14] 88.9 89.2 91.0 91.5 94.5 94.1 95.7 95.9
InfoGCN [17] 89.4 90.6 91.3 92.3 95.2 95.4 96.2 96.7
HD-GCN [18] - - 92.4 93.0 - - 96.6 97.0
DD-GCN [19] - 90.5 - 92.6 - 95.7 - 96.9
SelfGCN [23] - - - 93.1 - - - 96.6
STC-Net[22] - - 92.5 93.0 - - 96.7 97.1
HCN [7] - - 84.3* - - - 89.9* -
HCN w/CSRE - - $\text{85.1}^{\uparrow{0.8}}$ $\text{91.2}^{\uparrow{1.3}}$
2S-AGCN [4] 88.9 89.2 91.0 91.5 94.5 94.1 95.7 95.9
2S-AGCN w/CSRE $\text{89.4}^{\uparrow{0.5}}$ $\text{89.8}^{\uparrow{0.6}}$ $\text{91.5}^{\uparrow{0.5}}$ $\text{91.9}^{\uparrow{0.4}}$ $\text{94.7}^{\uparrow{0.2}}$ $\text{94.6}^{\uparrow{0.5}}$ $\text{96.1}^{\uparrow{0.4}}$ $\text{96.3}^{\uparrow{0.4}}$
CTR-GCN [5] 89.8 90.2 92.0 92.4 94.8 94.8 96.3 96.8
[…]
Hyperformer [10] 90.3 91.1 92.0 92.7 94.5 94.4 95.5 96.2
[…]
BlockGCN [11] 90.9 - - 93.1 95.4 - - 97.0
BlockGCN w/CSRE $\textbf{91.3}^{\uparrow{0.4}}$ $\underline{91.2}$ 92.6 $\textbf{93.3}^{\uparrow{0.2}}$ $\textbf{95.6}^{\uparrow{0.2}}$ 95.4 96.7 $\textbf{97.2}^{\uparrow{0.2}}$
NTU RGB+D 120. NTU120 dataset [37] adds 57367 new skeleton sequences and 60 new action classes to the original NTU60 dataset. There are 32 various configurations in it, each of which depicts a different location and background. The authors offered the cross-subject (X-sub) and cross-setup (X-set) as two benchmark evaluations. In the X-sub, sequences from 53 subjects are for training, and the other 53 subjects are for testing. In the X-set, skeleton sequences are split by setup ID. Samples from even set-up IDs are used for training, and the odds are used for evaluation.
Northwestern-UCLA. NW-UCLA dataset [38] contains 1494 video clips of 10 different actions captured from three Kinect cameras. We follow the same evaluation protocol in [38]: the first two cameras for training and the other for testing.
[…]
Dataset NTU 120
Setting X-Sub X-Set
Method/Streams J B 2S 4S J B 2S 4S
Shift-GCN [36] 80.9 - 85.3 85.9 83.2 - 86.6 87.6
TA-CNN [14] 84.0 85.1 87.8 88.2 85.3 86.3 89.0 89.6
InfoGCN [17] 84.2 86.9 88.2 89.2 86.3 88.5 89.4 90.7
HD-GCN [18] - - 89.1 89.7 - - 90.6 91.2
DD-GCN [19] - 86.1 - 88.9 - 87.6 - 90.2
SelfGCN [23] - - - 89.4 - - - 90.1
STC-Net [22] - - 89.3 89.9 - - 90.7 91.3
2S-AGCN [4] 83.8 84.9 87.7 88.1 85.3 86.3 89.0 89.6
2S-AGCN w/CSRE $\text{84.1}^{\uparrow{0.3}}$ $\text{85.1}^{\uparrow{0.2}}$ $\text{87.9}^{\uparrow{0.2}}$ $\text{88.4}^{\uparrow{0.3}}$ $\text{86.0}^{\uparrow{0.7}}$ $\text{86.6}^{\uparrow{0.3}}$ $\text{89.5}^{\uparrow{0.5}}$ $\text{90.0}^{\uparrow{0.4}}$
CTR-GCN [5] 84.9 85.7 88.7 88.9 86.7 87.5 90.1 90.5
CTR-GCN w/CSRE $\text{85.6}^{\uparrow{0.7}}$ $\text{86.7}^{\uparrow{1.0}}$ $\text{89.3}^{\uparrow{0.6}}$ $\text{89.5}^{\uparrow{0.6}}$ $\text{87.1}^{\uparrow{0.4}}$ $\text{88.2}^{\uparrow{0.7}}$ $\text{90.6}^{\uparrow{0.5}}$ $\text{91.0}^{\uparrow{0.5}}$
Hyperformer [10] 86.1 87.4 88.9 89.9 87.8 89.0 90.6 91.2
Hyperformer w/CSRE $\text{86.8}^{\uparrow{0.7}}$ $\underline{\text{88.1}}^{\uparrow{0.7}}$ $\underline{\text{89.4}}^{\uparrow{0.5}}$ $\underline{\text{90.3}}^{\uparrow{0.4}}$ $\underline{\text{88.3}}^{\uparrow{0.5}}$ $\underline{\text{89.2}}^{\uparrow{0.2}}$ $\textbf{91.0}^{\uparrow{0.4}}$ $\underline{\text{91.6}}^{\uparrow{0.4}}$
BlockGCN [11] 86.9 - - 90.3 88.2 - - 91.5
BlockGCN w/CSRE $\textbf{87.2}^{\uparrow{0.3}}$ 88.4 89.6 $\textbf{90.5}^{\uparrow{0.2}}$ $\textbf{88.7}^{\uparrow{0.5}}$ 89.5 91.0 $\textbf{91.8}^{\uparrow{0.3}}$
4.2 Implementation Details
To fully validate the effectiveness and generalizability of CSRE, we take five different model-based approaches as baseline models. (1) CNN-based method. We select HCN [7] as the CNN-based method to validate CSRE. We reproduce the experimental results using the released code and conduct our CSRE on it. We follow the training recipes described in the paper. (2) GCN-based method. 2S-AGCN [4] and CTR-GCN [5], the widely used GCN-based model for skeleton-based action recognition. For CTR-GCN, we follow their training strategies. Particularly, for 2S-AGCN, we update the training strategies from CTR-GCN, which can improve the performance significantly. Moreover, to validate whether CSRE is still applicable to the latest models, we also experiment on the combination of proposed CSRE and BlockGCN [11]. (3) Transformer-based method. For transformer-based methods, we choose Hyperformer as the baseline.
@@ 4.2 Implementation Details
AGC-LSTM [39] - - 93.3* -
DC-GCN+ADG [40] - - 95.3 -
Shift-GCN [36] 92.5 - 94.2 94.6
HD-GCN [18] - - 96.6 96.9
DD-GCN [19] - - - 96.7
SelfGCN [23] - - - 96.8
STC-Net [22] - - 96.8 97.2
2S-AGCN [4] 92.0 92.2 95.0 95.5
2S-AGCN w/CSRE $\text{93.8}^{\uparrow{1.8}}$ $\text{93.1}^{\uparrow{0.9}}$ $\text{95.9}^{\uparrow{0.9}}$ $\underline{\text{97.0}}^{\uparrow{1.5}}$
CTR-GCN [5] 94.6 91.8 94.2 96.5
CTR-GCN w/CSRE $\text{94.8}^{\uparrow{0.2}}$ $\text{94.6}^{\uparrow{2.8}}$ $\text{95.7}^{\uparrow{1.5}}$ $\textbf{97.2}^{\uparrow{0.7}}$
InfoGCN [17] 93.8 94.2 95.5 96.1
Info-GCN w/CSRE $\text{94.2}^{\uparrow{0.4}}$ $\text{95.5}^{\uparrow{1.3}}$ $\text{95.9}^{\uparrow{0.4}}$ $\text{96.5}^{\uparrow{0.4}}$
Hyperformer [10] 92.7 95.0 95.0 96.6
Hyperformer w/CSRE $\text{95.3}^{\uparrow{2.6}}$ $\text{95.3}^{\uparrow{0.3}}$ $\text{96.3}^{\uparrow{1.3}}$ $\underline{\text{97.0}}^{\uparrow{0.4}}$
BlockGCN [11] 95.5 - - 96.9
BlockGCN w/CSRE $\textbf{95.9}^{\uparrow{0.4}}$ 95.8 97.0 $\textbf{97.2}^{\uparrow{0.3}}$
For the details of contrastive learning, the size of the established memory banks is the length of the dataset. The dimensions of the attentive features $\bm{t}$ and $\bm{s}$ is 256. For all datasets, the number of sampling instances $N^{+}_{H}$, $N^{-}_{H}$ and $N^{-}_{R}$ are set as 128, 512, and 512, respectively. To select the model with the best performance, we experiment with temperatures $\tau$ of 0.5, 0.8, and 1.0. All experiments are conducted with Pytorch [41] deep learning library on four GeForce GTX NVIDIA 3090 GPUs. Moreover, to avoid random noise and ensure experiment reproducibility, we adopt the same random seed as the baseline methods and fix it.
@@
CTR-GCN [5] 84.9
CTR-GCN w/GAP [32] 85.5
CTR-GCN w/SkeletonGCL [33] 85.6
CTR-GCN w/FR Head [34] 85.5
CTR-GCN w/CSRE 85.6
Hyperformer [10] 86.1
Hyperformer w/CSRE 86.8
4.3 Comparison With The State-of-The-Art
We compare our method with previous SOTA methods in Tables 1, 2 and 3. To fully validate the proposed CSRE, we report the experimental results in a single modality and the ensemble results of multi-modality. $J$ means the joint stream, $B$ denotes the bone stream, $2S$ denotes the ensemble results of the above two streams by default, and $4S$ incidents the ensemble results of additional joint motion and bone motion streams. Particularly, due to the design of their model architecture, the experimental results of VA-LSTM [42], ACG-LSTM [39] and HCN [7] are the ensemble results of two streams. Therefore, we classify the results into $2S$.
Stable Improvements. As shown in Tables 1, 2 and 3, combined with CSRE, all five various baselines achieve solid gains on these five benchmarks of the three datasets over different modalities. Take CTR-GCN on the Xsub benchmark of the NTU120 dataset as a baseline example; it improves by 0.7% (84.9% to 85.6%) in the joint stream, 1.0% (85.7% to 86.7%) in the bone stream.
Ensemble Improvements. As shown in Tables 1, 2 and 3, the single-modal performance improvements brought by CSRE can be extended to the results of multi-stream fusion (2S and 4S). It improves CTR-GCN by 0.6% (88.7% to 89.3%) in two streams fusion results and 0.6% (88.9% to 89.5%) in four streams fusion results on the Xsub benchmark of the NTU120.
General Applicability. It can be observed that the proposed CSRE framework is applicable to various baselines, including the recent BlockGCN [11]. The results of CSRE with BlockGCN surpass other experimental results and achieve state-of-the-art techniques.
4.4 Comparison With Other Contrastive Learning Methods
We compare our method with the recent contrastive learning methods in supervised skeleton-based action recognition. The comparison is conducted on the X-sub benchmark of the NTU120 dataset. As shown in Table 4, our CSRE achieves competitive improvements. When conducting experiments based on CTR-GCN [5], the proposed CSRE achieves the same or even higher improvements compared to other contrastive learning methods.
Furthermore, previous contrastive learning methods [32, 36, 34] are only applicable to some GCN-based models due to their design, which limits their generalizability. Specifically, GAP [32] relies on the splitting of spatial dimensions according to human body regions for fine-grained feature alignment, which needs to maintain the spatial dimension in the aggregations. FR-Head [34] is imposed on different stages of GCNs to build a multi-level refinement for stronger supervision, which depends on the design of the mainstream GCN-based method heavily and demonstrates weaker generalization. SkeletonGCL [33] suggests that due to the high-order structural information in graphs of the GCN blocks, graph contrast can better benefit graph convolution learning. Thus, SkeletonGCL only incorporates GCN-based methods.
To take advantage of our method, we combine the CSRE with Hyperformer [10], which is a transformer-based method. Our method outperforms previous contrastive learning methods by 1.2%, which is a promising result.
4.5 Ablation Study